OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2

OpenStudy (muzzack):

you mean \[x^2 + 5x = -2\]

OpenStudy (anonymous):

yes sir

OpenStudy (muzzack):

well x = \[\frac{ 1 }{ 2 }(\sqrt{17-5})\]

OpenStudy (muzzack):

or

OpenStudy (muzzack):

x = \[\frac{ 1 }{ 2 }(-5-\sqrt{17})\]

OpenStudy (anonymous):

thank you sir I appreciate you taking out your time to help

OpenStudy (jebonna):

ur swag

OpenStudy (anonymous):

?